Pardon Me, Inc., recently issued new securities to finance a new TV show. The project cost $40 million, and the company paid 1.9 million in flotation costs. In addition, the equity issued had a flotation cost of 7 percent of the amount raised, whereas the debt issued had a flotation cost of 3 percent of the amount raised. If the company issued new securities in the same proportion as its target capital structure, what is the company’s target debt-equity ratio?
Answer #1

Cost of Project = $40,000,000
Flotation Cost = $1,900,000

Total Cost = Cost of Project + Flotation Cost
Total Cost = $40,000,000 + $1,900,000
Total Cost = $41,900,000

Weighted Average Flotation Cost = Flotation Cost / Total Cost
Weighted Average Flotation Cost = $1,900,000 / $41,900,000
Weighted Average Flotation Cost = 0.04535

Let Weight of Debt be x and Weigh of Equity be (1 - x)

Weighted Average Flotation Cost = Weight of Debt * Flotation Cost of Debt + Weight of Equity * Flotation Cost of Equity
0.04535 = x * 0.03 + (1 - x) * 0.07
0.04535 = x * 0.03 + 0.07 - x * 0.07
x * 0.04 = 0.02465
x = 0.61625

Weight of Debt = 0.61625

Weight of Equity = 1 - 0.61625
Weight of Equity = 0.38375

Debt-Equity Ratio = Weight of Debt / Weight of Equity
Debt-Equity Ratio = 0.61625 / 0.38375
Debt-Equity Ratio = 1.61